OpenStudy (anonymous):

the polynomial function with roots 1, –2, and 5

OpenStudy (saifoo.khan):

x = 1 (x-1)=0 x = -2 (x+2)=0 x = 5 (x-5) =0 Finally, (x-1)(x+2)(x-5) = 0 Solve.

OpenStudy (anonymous):

i need to slove this? (x-1)(x+2)(x-5) = 0

OpenStudy (saifoo.khan):

Yes Sir.

OpenStudy (anonymous):

x=1, -2, 5 now what?

OpenStudy (saifoo.khan):

Multiply it out.

OpenStudy (anonymous):

what do u mean by multiply it out?

OpenStudy (saifoo.khan):